In 2011, skateboarding legend Micheal Jordan's age was twice that of Olympic Star Lebron James minus 0 years. The sum of their ages in 2011 was 72. How old were Micheal and Lebron in 2011?
Answer by stanbon(75887)

In 2011, skateboarding legend Micheal Jordan's age was twice that of Olympic Star Lebron James minus 0 years. The sum of their ages in 2011 was 72. How old were Micheal and Lebron in 2011?
-----------------
Equations:
m = 2j
m + j = 72
-------------------
Substitute for "m" and solve for "j":
2j + j = 72
3j = 72
j = 24 (Lebron's age in 2011)
---
Solve for "m":
m + j = 72
m + 24 = 72
------------
m = 48 (Jordan's age in 2011)
---------------
Cheers,
Stan H.
===============